OpenStudy (anonymous):

2log(x+3)=2x+3 solve the equation and round to the nearest hundredth

OpenStudy (anonymous):

definitely a calculator question... you're basically solving: x+3 = 10^(x+3/2) do you have a graphing calculator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok. Here's one way to do it. Solve for 0 so you'll have 2log(x+3) - 2x -3 = 0. Now stick this left side expression as a function in your calculator and graph it. Most calculators can locate zeros of functions (or x-intercepts). Let your calculator take it from there.

OpenStudy (anonymous):

i forgot how to calculate the zeros of functions

OpenStudy (anonymous):

I can only refer to the calculator that I use (TI-83). I'm not familiar with Casios or HPs... What do you have?

OpenStudy (anonymous):

TI-84 plus

OpenStudy (anonymous):

maybe you could use an online graphing calculator...

OpenStudy (anonymous):

its tellin me its somewhere around -1.25

OpenStudy (anonymous):

perfect... so, after you stick in the function into y1, press 2nd trace to get to the calc menu. In the calc menu, you'll have a choice for zero (or x-intercept) of a function. It will then ask you if the function you want to locate the zeros for is displayed on the screen. If it is press enter. Then it will ask you for left/right bounds. You'll have to use the cursor keys to move the curser onscreen so that the cursor is left of the x-intercept (left bound) and then right of the x-intercept(right bound). When it asks you for a guess just press enter. After a while, the calculator will give you it's answer.

OpenStudy (anonymous):

thanks for all the help i got it

OpenStudy (anonymous):

To check if your answer is correct, stick that value back into your original equation..

OpenStudy (anonymous):

np